# Introducing Spotlight: A Novel Approach for Generating Captivating Key Information from Documents

### v1

@@ Abstract
In this paper, we introduce Spotlight, a novel paradigm for information extraction that produces concise, engaging narratives by highlighting the most compelling aspects of a document. Unlike traditional summaries, which prioritize comprehensive coverage, spotlights selectively emphasize intriguing content to foster deeper reader engagement with the source material. We formally differentiate spotlights from related constructs and support our analysis with a detailed benchmarking study using new datasets curated for this work. To generate high-quality spotlights, we propose a two-stage approach: fine-tuning a large language model on our benchmark data, followed by alignment via Direct Preference Optimization (DPO). Our comprehensive evaluation demonstrates that the resulting model not only identifies key elements with precision but also enhances readability and boosts the engagement value of the original document.
@@
1Ankan Mullick   1Sombit Bose   1Rounak Saha   2Ayan Kumar Bhowmick 2Aditya Vempaty    2Prasenjit Dey    2Ravi Kokku    1Pawan Goyal    1 Niloy Ganguly {ankanm, sbcs.sombit, runk20}@kgpian.iitkgp.ac.in {pawang, niloy}@cse.iitkgp.ac.in, {ayan, aditya, prasenjit, ravi}@emergence.ai 1Computer Science Engineering, IIT Kharagpur, India. 2Emergence AI
@@
Table 1: Qualitative Analysis of different metrics to evaluate and characterize various methods for deriving information from a source document.
@@ 1 Introduction
Unlike traditional formats, spotlights offer the narrative depth of summaries while uniquely aiming to generate interest in the full text. [^1] Note that even the summarization can be target-specific, some examples include query focused summarization Vig et al. (2022); Baumel et al. (2016); Liu et al. (2024), layman summarization Goldsack et al. (2023); Salaün et al. (2022); Luo et al. (2022), goal oriented Goldstein et al. (2007); Ham et al. (2020); Jin et al. (2024), etc. summaries with controllable readability He et al. (2020); Zhang et al. (2023); Ribeiro et al. (2023), summarizing documents with respect to key elements Ryu et al. (2024); Wang et al. (2023). However, the objective of these methods is generally to achieve maximum coverage and deliver a quick, faithful read—potentially sparing the reader from having to engage with the full main text, which appears to be in contrast to the objective of Spotlight. Moreover, the concept of Spotlights can also be extended to cater to these objectives, e.g., query-focused spotlight. They strike a careful balance between readability, engagement, and fidelity, functioning as standalone narratives. Much like an author’s spotlight talk at a conference, they are concise yet compelling, designed to draw readers in and encourage deeper exploration. While these five condensation techniques may exist along a continuum, they possess distinct characteristics that justify separate classification. Table 1 outlines their comparative features, such as source fidelity, level of intrigue, abstractiveness, conciseness, narrative coherence, readability, and the distribution of content coverage.
While extensive research exists on summarization and highlight extraction (see Appendix E), spotlight generation remains unexplored. In this work, we introduce the concept, quantitatively differentiate spotlights from summaries, and propose effective generation methods. Although no dedicated spotlight datasets currently exist, we identify and adapt existing sources—such as news articles, CSPubSum, Wikipedia, and conference presentations—to align with spotlight characteristics. These are compared with corresponding summaries based on the framework defined through Table 1.
@@ 2.1 Spotlight Datasets
Research Presentation: We curate this dataset by collecting 1,230 audio transcripts of NeurIPS 2022  [^2] https://nips.cc/virtual/2022/events/Spotlight and 2023  [^3] https://nips.cc/virtual/2023/events/spotlight-posters-2023 Spotlight Poster presentations. These audio transcripts are derived from 5-minute self-narrative videos, which focus on presenting the most intriguing and important aspects of the research papers. This concise format is meant to capture the attention of the audience and encourage them to explore the full paper. Since these audio transcripts are generated from automated transcription services, they often contain errors that could hinder readability and interpretation. Therefore, we apply transcript cleaning techniques like contraction expansion, punctuation removal, etc. to improve the quality and utility of these transcripts.
@@ 2.2 Summary Generation
We generate summaries by human as well as using LLMs.
@@ 3 Spotlight: Characterization
Table 3: Feature level analysis of Spotlight(Spot) vs. Summary(Summ), Entropy, Compactness, Flesch Kincaid and Gunning score lower values signify better, whereas for Skew50, Flesch Reading score and Interest-Generation higher values signify a better result as discussed in Section 3, Additional Results in Appendix Table 10 and 11
@@ 3.1 Spotlight-Summary: Distinct Features
We do a comparitive study on four datasets for 60 annotated samples (each instance has a document, corresponding spotlight and manual summary).
@@
(B) Information Compactness: Information compactness refers to the degree to which a condensation distills essential details of the original content. We use the average IDF [^4] IDF stands for Inverse Document Frequency. IDF is a standard metric to measure the informativeness of a word. score to measure compactness. We calculate the IDF scores for each word of the condensation and then do the average. From Table 3 it is clearly seen that the average IDF value is always higher for summary than spotlight indicating more information overload (compactness) in summary than spotlight.
@@
Figure 1: Comparison between a summary and a spotlight generated from the same document.(interest generation phrases are highlighted in bold and underlined)
(F) Interest Generation: We conducted a human annotation study focused on interest generation. For each dataset, 50 documents with manually annotated summaries were selected, and two independent annotators were asked to indicate whether the summary or the spotlight was more likely to generate interest for deeper exploration among readers. Annotators also had the option to state that both formats generate similar levels of interest. The overall inter-annotator agreement was 0.87. As shown in Column F of Table 3, annotators overwhelmingly rated spotlights generate more interest than summaries.We find that the score differences (spotlight vs summary) are statistically significant ($p<0.05$) as per McNemar’s Test.
@@ 4 Spotlight Generation Process
Open Source LLMs: We use various small-size large language models (LLMs) - Mistral-7b Jiang et al. (2023) and Llama [^5] https://ai.meta.com/llama/ based - Llama2-7b and Llama3-8b models. We experiment with Mistral-7b (abbreviated as Mst-7b-FT), Llama2-7b and Llama3-8b fine tuning (abbreviated as L2-7b-FT and L3-8b-FT respectively), general instruction tuned vanilla models (Mst-7b-VA, L2-7b-VA and L3-8b-VA respectively) and corresponding DPO models. We report the best performing Llama3-8b LLM-based outcomes in Table 5 and other details in the Appendix Tables 20, 22, 24 and 26.
@@ 5.1 Baseline Approaches
(B) Large-Size LLMs: We apply example-based prompting (1-shot) on large-sized LLMs like ChatGPT [^6] https://chat.openai.com/, GPT-4777https://openai.com/gpt-4 and Gemini-Pro [^8] https://gemini.google.com/ to check their capability to generate good quality spotlights. Due to space shortage, ChatGPT, and Gemini-Pro evaluations are in Appendix Tables 20, 22, 24 and 26.
@@ 5.2 Evaluation
(A) Traditional: We check the competence of different models with traditional evaluations metrics like (i) Rouge1 (R1), Rouge2 (R2) and  Lin (2004), (ii) BERT-F1 (BeF1) Zhang et al. (2019).
@@
(C) GPT4 Critique Mullick et al. (2024): GPT-4 is asked to rate the spotlight – whether it is good or bad with respect to the ground truth spotlight and whether it generates interest to read the source document.
@@ 5.3 Experimental Settings
We use 80GB A100 GPU, 210MHz clock cycle with Huggingface TRL and PEFT libraries built on top of Huggingface Transformers to conduct the experiments of fine-tuning and alignment of pretrained models and for their evaluation purpose. Additionally, we use the NLTK, spaCy, and NumPy packages for further analysis and ablations. Code details are in Link.
@@ 6 Experimental Results and Discussion
Table 4: Traditional, UniEval and GPT-4 Critique evaluations on CSPubSum Dataset - comparison of baselines with SFT, Additional Results in Appendix Table 19, 21, 23 and 25
@@
(A) Performance of Traditional Baseline Approaches: We explore different traditional approaches (LongFormer, T5-Fine-tune, BART-large, Pegasus, Falcon, TLDR) for spotlight generation and results for CSPubSum are shown in Table 4. Detailed results for the datasets are in Appendix Table - 19 (for News), 21 (for CSPubSum), 23 (for Wikipedia) and 25 (for Research Presentation) It is seen that even simple SFT-based approaches on LLMs perform much better than these traditional approaches across all metrics. This performance gap may be due to the advanced architectures and targeted optimizations of LLM models which are fine-tuned specifically for task performance and user preferences, whereas the older models do not benefit from the same level of refinement and cutting-edge techniques, leading to their comparatively lower performance.
@@
Table 5: Traditional, UniEval and GPT-4 Critique evaluations for best performing Llama3-8b and GPT-4 approach in % (best outcome in Bold) for each dataset, Additional evaluation details including other LLMs are in Appendix Table 20, 22, 24 and 26
@@
Table 6: Spotlight Characteristics Evaluation on Model Generated Spotlight, Entropy, Compactness, Flesch Kincaid and Gunning score lower values signify better result, but for Skew50 and Flesch Reading score higher values signify a better result as discussed in Section 3
@@
  • We explore different DPO configurations for utilization of data and fine-tuning approach (details in Appendix I.4) and the best result is obtained in our 2-step approach when half of the data is used for SFT followed by DPO for rest half data.
@@
  • To validate the GPT-4 critique scores against robust human judgment, we conducted an additional human annotation study similar to Mullick et al. (2024) to evaluate the alignment between human judgment and automated assessments for 20 documents for each dataset, details in Appendix I.2. We obtain an overall average Point-biserial correlation coefficient of 0.751 between the GPT-4 critique scores and human judgement.
(C) DPO Conformance to Spotlight Features: We do an extensive evaluation of the model-generated spotlights of DPO pipeline with the fine-tuned version across various spotlight characteristics, as detailed in Table 6 which shows that - (i) DPO (L3-8b) consistently improves readability across all datasets, producing spotlights that are easier to understand than fine-tune approach; (ii) Entropy reductions and increase in Skew50 across all datasets imply DPO maintains skewness in term distribution than fine-tune. (iii) Compactness decreases in DPO, indicating less information overload and more readability. In conclusion, DPO produces more structured, readable, and well-balanced spotlights compared to fine-tuned approach.
@@
(D) Divergence of DPO-generated Spotlights from the gold standard Summary: As outlined in Section 4, our spotlight generation process consists of two steps. First, we fine-tune the base model (FT), and then we apply the DPO method to the fine-tuned model to align the output closer to a spotlight and further from a summary. While DPO produces superior spotlights (Table 5), a question remains: does the spotlight diverge from the summary characteristics, as per the DPO objective? To answer this, we compare model-generated spotlights with gold-standard summaries (GPT-4 generated) for three approaches: vanilla (L3-8b-VA), fine-tuning (L3-8b-FT Spot), and DPO (DPO (L3-8b) Spot), as presented in Table 7 for entire datasets. Our findings indicate that the spotlights generated using SFT are closer to the reference summaries than those produced by DPO, confirming the deviation. Nonetheless, both the fine-tuned and DPO outputs outperform the vanilla model, which prompts us to investigate whether the spotlight generation process inadvertently produces quality summaries.

### v3

@@ Abstract
Analyzing and processing vast amounts of textual data presents significant challenges in efficiently extracting key information. In this paper, we introduce ‘Spotlight’, a novel paradigm for information extraction that produces concise, engaging narratives by highlighting the most compelling aspects of a document. Unlike highlights (fragmented key points) and traditional summaries, which prioritize comprehensive coverage, spotlights selectively emphasize intriguing content to foster deeper reader engagement with the source material. We formally differentiate spotlights from related constructs and support our analysis with a detailed benchmarking study using new datasets curated for this work. To generate high-quality spotlights, we propose a two-stage approach: fine-tuning a large language model on our benchmark data, followed by alignment via Direct Preference Optimization (DPO). Our comprehensive evaluation demonstrates that the resulting model not only identifies key elements with precision but also enhances readability and boosts the engagement value of the original document. Datasets and code are available at https://github.com/ankan2/Spotlight-EMNLP2025.
@@
1∗Ankan Mullick    1∗Sombit Bose    1Rounak Saha    2Ayan Kumar Bhowmick 2Aditya Vempaty    2Prasenjit Dey    2Ravi Kokku    1Pawan Goyal    1 Niloy Ganguly {ankanm, sbcs.sombit.24, runk20}@kgpian.iitkgp.ac.in {pawang, niloy}@cse.iitkgp.ac.in, {ayan, aditya, prasenjit, ravi}@emergence.ai 1IIT Kharagpur, India. 2Emergence AI
**footnotetext: Equal Contribution.
@@
Table 1: Qualitative Analysis of different metrics to evaluate and characterize various methods (Summary, Teaser, Headline, Highlight, and Spotlight) for deriving information from a source document.
@@ 1 Introduction
This paper introduces the concept of spotlights as an innovative step forward from traditional summaries and highlights. A spotlight is a concise, self-contained narrative, typically prepared by the document’s author, that highlights its most significant and engaging aspects. Unlike traditional formats, spotlights offer the narrative depth of summaries while uniquely aiming to generate interest in the full text.***Note that even the summarization can be target-specific, some examples include query focused summarization Vig et al. (2022); Baumel et al. (2016); Liu et al. (2024), layman summarization Goldsack et al. (2023); Salaün et al. (2022); Luo et al. (2022), personal-based summarization Mullick et al. (2024c), goal oriented Goldstein et al. (2007); Ham et al. (2020); Jin et al. (2024), aspect-based summarization Mullick et al. (2024d) etc. summaries with controllable readability He et al. (2020); Zhang et al. (2023); Ribeiro et al. (2023), summarizing documents with respect to key elements Ryu et al. (2024); Wang et al. (2023). However, the objective of these methods is generally to achieve maximum coverage and deliver a quick, faithful read—potentially sparing the reader from having to engage with the full main text, which appears to be in contrast to the objective of spotlight. Moreover, the concept of spotlights can also be extended to cater to these objectives, e.g., query-focused spotlight, persona-based spotlight etc. They strike a careful balance between readability, engagement, and fidelity, functioning as standalone narratives. Much like an author’s spotlight talk at a conference, they are concise yet compelling, designed to draw readers in and encourage deeper exploration. While these five condensation techniques may exist along a continuum, they possess distinct characteristics that justify separate classification. Table 1 outlines their comparative features, such as source fidelity, level of intrigue, abstractiveness, conciseness, narrative coherence, readability, and the distribution of content coverage.
While extensive research exists on summarization and highlight extraction (see Appendix E), spotlight generation remains unexplored. In this work, we introduce the concept, quantitatively differentiate spotlights from summaries, and propose effective generation methods. Although no dedicated spotlight datasets currently exist, we identify and adapt existing sources—such as news articles, CSPubSum, Wikipedia, and conference presentations—to align with spotlight characteristics. These are compared with corresponding similar concepts (summary, teaser, headline, highlight based on the framework defined through Table 1.
@@ 2.1 Spotlight Datasets
Research Presentation: We curate this dataset by collecting 1,230 audio transcripts of NeurIPS 2022 ***https://nips.cc/virtual/2022/events/Spotlight and 2023 ***https://nips.cc/virtual/2023/events/spotlight-posters-2023 spotlight Poster presentations. These audio transcripts are derived from 5-minute self-narrative videos, which focus on presenting the most intriguing and important aspects of the research papers. This concise format is meant to capture the attention of the audience and encourage them to explore the full paper. Since these audio transcripts are generated from automated transcription services, they often contain errors that could hinder readability and interpretation. Therefore, we apply transcript cleaning techniques like contraction expansion, punctuation removal, etc. to improve the quality and utility of these transcripts.
@@ 2.2 Summary Generation
We generate summaries manually by annotators as well as by using LLMs.
@@ 3 Spotlight: Characterization
Table 3: Feature level analysis of Spotlight (Spot) vs. Summary (Summ), Entropy, Compactness, Flesch Kincaid and Gunning score lower values signify better, whereas for Skew50, Flesch Reading score and Interest-Generation higher values signify a better result as discussed in Section 3, Additional Results in Appendix Table 11 and 12
@@ 3.1 Spotlight-Summary: Distinct Features
We do a comparative study on four datasets for 60 annotated samples (each instance has a document, corresponding spotlight and manual summary).
@@
(B) Information Compactness: Information compactness refers to the degree to which a condensation distills essential details of the original content. We use the average IDF***IDF stands for Inverse Document Frequency. IDF is a standard metric to measure the informativeness of a word. score to measure compactness. We calculate the IDF scores for each word of the condensation and then do the average. From Table 3 it is clearly seen that the average IDF value is always higher for summary than spotlight indicating more information overload (compactness) in summary than spotlight.
@@
Figure 1: Comparison between a summary and a spotlight generated from the same document. The summary presents a concise, technical overview of the algorithm, detailing how adaptive vertex relocation minimizes approximation error with rigorous experimental validation. The spotlight describes a captivating mini-story, emphasizing the innovative breakthrough of liberating vertices from the contour to create elegant, engaging shape approximations. It minimizes emphasis on technical methodology and reduces the use of complex terminology, while effectively conveying the significance of the work in a way that encourages readers to explore the main article. Each phrase introduces a novel concept or method, prompting readers to wonder how exactly it works and why it improves results, thereby triggering a sense of deeper exploration by raising implicit questions in the reader’s mind.
The “interest generation phrases” are highlighted in bold and underlined).
(F) Interest Generation: We conducted a human annotation study focused on interest generation. For each dataset, 50 documents with manually annotated summaries were selected, and two independent annotators were asked to indicate whether the summary or the spotlight was more likely to generate interest for deeper exploration among readers. Annotators also had the option to state that both formats generate similar levels of interest. The overall inter-annotator agreement was 0.87. As shown in Column F of Table 3, annotators overwhelmingly rated spotlights as generating more interest than summaries. We find that the score differences (spotlight vs summary) are statistically significant ($p<0.05$) as per McNemar’s Test.
@@ 3.2 Spotlight-Summary: Similar Features
3.3 Spotlight-Highlight: Distinct Features
To explore the distinctness between spotlight and highlights, we have performed a similar mini-story study to Section 3.2. Details in Appendix Section G.3. In our analysis of distinct features (as described in Section 3.2), we observed that spotlights consistently outperform highlights in conveying a mini-story narrative. Annotator evaluations indicate that spotlights were rated as having a more engaging narrative as shown in Table 4. While both spotlights and highlights aim to distill key information from source documents, spotlights uniquely present a self-contained narrative that not only summarizes but also sparks curiosity and encourages further exploration. This narrative element is less pronounced in traditional highlights, which tend to list isolated key points rather than weaving them into a cohesive mini-story.
Dataset Spotlight(%) Highlights(%) News 80 20 CSPubSum 72 28 Wikipedia 72 28 Research Presentation 84 16
Table 4: Mini-story validation for spotlight vs. highlight analysis across all datasets
@@ 4 Spotlight Generation Process
Open Source LLMs: We use various small-size large language models (LLMs) - Mistral-7b Jiang et al. (2023) and Llama***https://ai.meta.com/llama/ based - Llama2-7b and Llama3-8b models. We experiment with Mistral-7b (abbreviated as Mst-7b-FT), Llama2-7b and Llama3-8b fine tuning (abbreviated as L2-7b-FT and L3-8b-FT respectively), general instruction tuned vanilla models (Mst-7b-VA, L2-7b-VA and L3-8b-VA respectively) and corresponding DPO models. We report the best performing Llama3-8b LLM-based outcomes in Table 6 and other details in the Appendix Tables 21, 23, 25 and 27.
@@ 5.1 Baseline Approaches
(B) Large-Size LLMs: We apply example-based prompting (1-shot) on large-sized LLMs like ChatGPT***https://chat.openai.com/, GPT-4***https://openai.com/gpt-4 and Gemini-Pro***https://gemini.google.com/ to check their capability to generate good quality spotlights. Due to space shortage, ChatGPT, and Gemini-Pro evaluations are in Appendix Tables 21, 23, 25 and 27.
@@ 5.2 Evaluation
(A) Traditional: We check the competence of different models with traditional evaluations metrics like (i) Rouge1 (R1), Rouge2 (R2),  Lin (2004) and (ii) BERT-F1 (BeF1) Zhang et al. (2019).
@@
(C) GPT4 Critique Mullick et al. (2024c): GPT-4 is asked to rate the spotlight – whether it is good or bad with respect to the ground truth spotlight and whether it generates interest to read the source document.
@@ 5.3 Experimental Settings
We use 80GB A100 GPU, 210MHz clock cycle with Huggingface TRL and PEFT libraries built on top of Huggingface Transformers to conduct the experiments of fine-tuning and alignment of pretrained models and for their evaluation purpose. Additionally, we use the NLTK, spaCy, and NumPy packages for further analysis and ablations. Further details are in Appendix F.
@@ 6 Experimental Results and Discussion
Table 5: Traditional, UniEval and GPT-4 Critique evaluations on CSPubSum Dataset - comparison of baselines with SFT, Additional Results in Appendix Table 20, 22, 24 and 26
@@
(A) Performance of Traditional Baseline Approaches: We explore different traditional approaches (LongFormer, T5-Fine-tune, BART-large, Pegasus, Falcon, TLDR) for spotlight generation and results for CSPubSum are shown in Table 5. Table 5 shows that traditional summarization models such as LongFormer, T5, BART, Pegasus, and Falcon struggle to generate effective spotlights, with lower ROUGE, BERT-F1, and UniEval scores. In contrast, even simple supervised fine-tuning of LLMs (Llama2, Mistral, Llama3) yields large improvements, confirming that LLMs can adapt well to the spotlight objective. This gap underscores that generating spotlights demands richer semantic modeling than what traditional summarization frameworks provide. Detailed results for the datasets are in Appendix Table - 20 (for News), 22 (for CSPubSum), 24 (for Wikipedia) and 26 (for Research Presentation) It is seen that even simple SFT-based approaches on LLMs perform much better than these traditional approaches across all metrics. This performance gap may be due to the advanced architectures and targeted optimizations of LLM models which are fine-tuned specifically for task performance and user preferences, whereas the older models do not benefit from the same level of refinement and cutting-edge techniques, leading to their comparatively lower performance.
@@
Table 6: Traditional, UniEval and GPT-4 Critique evaluations for best performing Llama3-8b and GPT-4 approach in % (best outcome in Bold) for each dataset, Additional evaluation details including other LLMs are in Appendix Table 21, 23, 25 and 27
@@
Table 7: Spotlight Characteristics Evaluation on Model Generated spotlight - Entropy, Compactness, Flesch Kincaid and Gunning score lower values signify better result, but for Skew50 and Flesch Reading score higher values signify a better result as discussed in Section 3
@@
  • We explore different DPO configurations for utilization of data and fine-tuning approach (details in Appendix I.4) and the best result is obtained in our 2-step approach when half of the data is used for SFT followed by DPO for the remaining half of the data.
@@
  • To validate the GPT-4 critique scores against robust human judgment, we conducted an additional human annotation study similar to Mullick et al. (2024c) to evaluate the alignment between human judgment and automated assessments for 20 documents for each dataset, details in Appendix I.2. We obtain an overall average Point-biserial correlation coefficient of 0.751 between the GPT-4 critique scores and human judgment.
Across the datasets, DPO applied on Llama3-8b consistently outperforms both vanilla and fine-tuned versions. While GPT-4 (1-shot) is strong in certain cases, particularly in UniEval, the DPO-trained Llama3-8b achieves the best balance across all metrics. This demonstrates that direct preference optimization is especially effective for spotlight generation.
(C) DPO Conformance to Spotlight Features: We do an extensive evaluation of the model-generated spotlights of DPO pipeline with the fine-tuned version across various spotlight characteristics, as detailed in Table 7 which shows that (i) DPO (L3-8b) consistently improves readability across all datasets, producing spotlights that are easier to understand than fine-tune approach; (ii) Entropy reductions and increase in Skew50 across all datasets imply DPO maintains skewness in term distribution than fine-tune. (iii) Compactness decreases in DPO, indicating less information overload and more readability. In conclusion, DPO produces more structured, readable, and well-balanced spotlights compared to fine-tuned approach. These observations confirm that DPO yields spotlights that are not only more structured and readable but also more faithful to the intended characteristics of spotlight generation, in contrast to the less targeted outputs of a purely fine-tuned approach.
@@
(D) Divergence of DPO-generated Spotlights from the gold standard Summary: As outlined in Section 4, our spotlight generation process consists of two steps. First, we fine-tune the base model (FT), and then we apply the DPO method to the fine-tuned model to align the output closer to a spotlight and further from a summary. While DPO produces superior spotlights (Table 6), a question remains: does the spotlight diverge from the summary characteristics, as per the DPO objective? To answer this, we compare model-generated spotlights with gold-standard summaries (GPT-4 generated) for three approaches: vanilla (L3-8b-VA), fine-tuning (L3-8b-FT Spot), and DPO (DPO (L3-8b) Spot), as presented in Table 8 for the entire datasets. Our findings indicate that the spotlights generated using SFT are closer to the reference summaries than those produced by DPO, confirming the deviation. Table  8 examines whether spotlight-trained models drift toward summary characteristics. The results indicate that spotlights produced by SFT are closer to gold summaries, whereas DPO outputs diverge more strongly, aligning with the intended spotlight objective. At the same time, when models are trained explicitly on summaries, they achieve higher traditional metrics but lose the engagement qualities of spotlights. This confirms that while spotlights and summaries overlap, optimizing for one does not automatically guarantee the qualities of the other. Nonetheless, both the fine-tuned and DPO outputs outperform the vanilla model, which prompts us to investigate whether the spotlight generation process inadvertently produces quality summaries.